14. The condensed electron configuration of silicon, element 14, is a. He]2s42p6 b. [Ne12p10 c. [Ne]3s23p2 d.[He12s4 e [He12s62p2 15. Enthalpy is an expression for the a. heat conten b. c. d. activation energy energy state reaction rate Ir Alíend hernmlo kcal/mol, it indicates the reaction is 16. If ?//reaction, is?100 kcal/mol, it indicates the reaction is a. endothermic b. nstable c. d. in need of a catalyst exothermic 17. Which of the following represent the Lewis structure for CI? :Cl: :Cl: c. Cl d. .Cl: 3 CHM2045 Pre-Post Test
media%2Ff10%2Ff1031a53-91ff-49af-8217-f5
Answer #1

14. expression for Si,

c.

--

15. enthalpy is,

a.

--

16. dHrxn = -ve means,

d.

--

17. lewis structure for Cl,

b.

--

2. using,

P1V1 = P2V2

(i) Volume (V2) = 1.5 atm x 554 ml/2.3 atm = 361.30 ml

(ii) Pressure (P2) = 1.5 atm x 554 ml/225 ml = 3.69 atm

(iii) Volume (V2) = 1.5 atm x 554 ml/(700/760) atm = 902.23 ml

(iv) Pressure (P2) = 1.5 atm x 554 ml/(1.2 L x 1000) ml = 0.6925 atm

--

3. for He in a cyllinder

using,

V1/T1 = V2/T2

(i) Volume (V2) = 5.5 L x 85 oC/35 oC = 13.36 L

(ii) Temperature (T2) = 8.2 L x 35 oC/5.5 L = 52.20 oC

(iii) Volume (V2) = 5.5 L x 5 oC x 1000/35 oC = 785.71 ml